OpenStudy (anonymous):

How to solve compound inequalities? I have been stuck on these for awhile now and have watched a few videos but still dont understand.. |x|>9 |b+6|<(with a line under it)5 |y+6|>13 |11-2b|-6>11 I dont want answers for all, but would love for someone to explain so i can understand and do it on my own.

OpenStudy (anonymous):

So are those 1 whole problem or 4 separate problems?

OpenStudy (jdoe0001):

\(\bf |x|>9\implies \begin{cases} +(x)>9\\ \quad \\ \bf -(x)>9 \end{cases}\\\\ \quad \\ |b+6|\le 5\implies \begin{cases} +(b+6)\le 5\\ \quad \\ \bf -(b+6)\le 5 \end{cases}\\\\ \quad \\ |y+6|>13\implies \begin{cases} +(y+6)>13\\ \quad \\ \bf -(y+6)>13 \end{cases}\\\\ \quad \\ |11-2b|-6>11\implies |11-2b| > 11+6\implies |11-2b| > 17\\ \quad \\ \bf |11-2b| > 17\implies \begin{cases} +(11-2b)>17\\ \quad \\ \bf -(11-2b)>17 \end{cases}\) each absolute value expression, will give you 2 values for "x", 1 for each scenario